# Fluid dynamics, Statistical physics, Theoretical physics

### v2

@@ 1 Introduction
In his 1964 Lectures [Feynman et al.(2011)Feynman, Leighton, and Sands], Richard Feynman identified the analysis of turbulent fluids as a problem "left over from a long time ago… Nobody in physics has really been able to analyze it mathematically satisfactorily." Decades later, this challenge, along with the related problem of non-perturbative dynamics in Yang-Mills gauge theory, has remained a frontier of theoretical science. These phenomena are governed by strongly nonlinear equations that defy standard analytical techniques, forcing a reliance on phenomenological models or numerical simulation.
@@ A Brief Historical Context
The challenge of deriving a statistical theory of turbulence directly from the Navier-Stokes equations has a long history. Notably, Eberhard Hopf pioneered the use of functional methods, introducing an equation for the characteristic functional of the velocity field and envisioning a universal turbulent attractor governing the long-time statistics (see a recent review [Ohkitani(2020)]). However, Hopf’s functional equation proved intractable, leaving the nature of the attractor unclear for decades. Now, with the loop equation replacing the Hopf functional equation, we have its solution —the Euler ensemble —as the realization of Hopf’s conjecture about the turbulent attractor.
Later, the loop space formalism emerged, rooted in the Makeenko-Migdal (MM) equations of the late 1970s and early 1980s [Makeenko and Migdal(1979), Makeenko and Migdal(1981), Migdal(1983)]. These were immediately recognized in the high-energy physics community as the first exact, non-perturbative dynamical equations for QCD. However, the initial hope of finding an analytical solution for quark confinement soon faded. The mathematical formulation, while geometrically insightful, proved analytically intractable, hampering further progress and shifting the community’s focus toward lattice gauge theory. The loop approach offered a different path compared to Hopf’s formulation, focusing on circulation observables.
The scope of the formalism was extended to fluid dynamics in the early nineties with the derivation of the loop equations for the Navier-Stokes system [Migdal(1995b)]. Within this framework, a WKB solution was derived for the limit of large, smooth loops. This solution established a theoretical "area law" for the statistics of velocity circulation. A transformation to a dual "momentum loop space" was also shown to reduce the problem to a singular, one-dimensional equation. For decades, however, further progress was impeded by the mathematical challenges posed by this equation and by a lack of high-precision data with which to test the theory’s predictions.
The theoretical landscape evolved considerably after 2019, when high-resolution direct numerical simulations (DNS) by Sreenivasan and collaborators reported results consistent with the area law derived nearly three decades earlier [Iyer et al.(2021)Iyer, Bharadwaj, and Sreenivasan, Pumir(2021)]. This numerical evidence renewed interest in the approach and provided the impetus to resolve the long-standing mathematical challenges of the momentum loop equation. The complete analytical solution that emerged, realizing aspects of Hopf’s original vision via loop space, is the subject of this review. (See also [Migdal(2024h)] for further discussion on the connection to Hopf’s work).
@@
This theoretical structure, initially derived from physical principles, has recently received significant independent validation. From the number theory perspective, the statistical properties of the Euler ensemble at finite $N$ were studied in [Basak and Zaharescu(2025)] by combinatorial methods, confirming and extending the results and conjectures made in the first paper [Migdal(2023)]. In addition, the distribution of radii of regular star polygons with unit side lengths was recently reproduced by Debmalya Basak using rigorous methods of number theory (private communication, to be published). From a Navier-Stokes dynamics perspective, a rigorous analysis by Brue and De Lellis [Elia and Lellis(2025)] has confirmed that the Euler ensemble is an exact solution to the discrete version of the loop equation (Theorem 10.6 in their paper). While the final step—proving the existence of the continuum limit of the fluid dynamics observables as the number of loop segments $N\to\infty$—remains an open mathematical problem, their result provides a firm mathematical foundation for the core of the theory. It confirms that the Euler ensemble is not an approximation but an exact solution at the level of the regularized theory.
@@
This mathematical validation is now strongly complemented by new high-resolution 4K DNS results from Sreenivasan and collaborators [Sreenivasan and Rodhiya, Akash(2026)]. These results verify the theory’s quantitative predictions with high precision, representing a significant development. As discussed in Sections 4.1 and 8, the DNS data for key observables—such as the $E\propto t^{-5/4}$ decay law and the universal, nonlinear effective index—are in excellent, parameter-free agreement with the theoretical curves derived from the Euler ensemble (see Fig. 2). Together, these mathematical and numerical results provide strong support for the loop space calculus in the Navier-Stokes equation, confirming the Euler ensemble as a robust solution for decaying turbulence.
@@
This review is organized as follows. The discussion begins with an introduction to the general loop space approach, applicable to both abelian fields in fluid dynamics and nonabelian fields in Yang-Mills theory. The approach is then applied to a hierarchy of turbulence problems, beginning with the foundational case of decaying hydrodynamic turbulence, then the transport of passive scalars, and finally the more complex case of magnetohydrodynamic (MHD) turbulence. A subsequent section situates these analytical solutions within the field’s historical context and clarifies their relationship to classical paradigms of turbulence theory. A particularly striking consequence of this framework emerges in gauge theory. In the Yang–Mills gradient flow, the same loop-space formalism leads to an exact analytic solution—the Hodge-dual matrix surface—which solves the fixed-point Yang–Mills loop equation by harmonic map [Migdal(2025a)]. This result establishes a concrete geometric foundation for the QCD confining string and provides a direct analytical bridge between turbulence, gauge theory, and string dualities.
@@ 2 The Loop Space Approach: From Nonlinear PDEs to Linear Diffusion
Here, the gauge field $A_{\mu}$ belongs to some Lie algebra, $A_{\mu}=\mathrm{i}\sum T_{a}A^{a}_{\mu}$, which we do not need to specify. The equation for the trace of the ordered path integral
@@
The transformation of both equations to the loop space diffusion equation is based on the operator calculus invented by Feynman [Feynman(1951)]. As applied to our case of the loop functionals for fluid dynamics and the gauge theory, it allows us to write both of them as the path ordered exponential times the unit operator
@@
This fundamental geometric identity, proven in the [Migdal(2025b)] and our Appendix A, is identical in abelian and non-abelian cases, except that a group space trace is taken in the Yang-Mills case. In both cases, it reveals the geometric meaning of the loop functional as a parallel transport of the covariant derivative operator around a loop $C$, reducing to a c-number for a closed loop. The reduction of the ordered exponential to a c-number is also supported by the Magnus expansion [Magnus(1954)]. This distinction is subtle but critical. In much of the mathematical literature such as [Göckeler and Schücker(1987)], the Wilson loop (the L.H.S. of our identity) is taken as the definition of parallel transport. Our proof, however, establishes this correspondence as a rigorous equality derived from the more fundamental, anchored operator (the R.H.S.). This approach mirrors a historical pattern in theoretical physics where a concept, once viewed as a powerful analogy, is later shown to be a provable, first-principles identity.
@@ 3 Momentum loop equation
The relation between the loop operator and the discontinuity of the momentum loop $\Delta P_{\nu}(\theta)$ was discovered and investigated in earlier papers in QCD [Migdal(1986), Migdal(1995a), Migdal(1998)]. The Momentum loop equation for QCD involves contact terms that provide boundary conditions in loop space for self-intersecting loops. We are not studying these terms in the present paper, restricting ourselves to the Yang-Mills gradient flow. For QCD, this equation corresponds to the WKB limit of large loops without self-intersections, when some form of the area law is expected as an asymptotic solution.

### v3

@@ 1 Introduction
In his 1964 Lectures [Feynman], Richard Feynman identified the analysis of turbulent fluids as a problem "left over from a long time ago… Nobody in physics has really been able to analyze it mathematically satisfactorily." Decades later, this challenge, along with the related problem of non-perturbative dynamics in Yang-Mills gauge theory, has remained a frontier of theoretical science. These phenomena are governed by strongly nonlinear equations that defy standard analytical techniques, forcing a reliance on phenomenological models or numerical simulation.
@@ A Brief Historical Context
The challenge of deriving a statistical theory of turbulence directly from the Navier-Stokes equations has a long history. Notably, Eberhard Hopf pioneered the use of functional methods, introducing an equation for the characteristic functional of the velocity field and envisioning a universal turbulent attractor governing the long-time statistics (see a recent review [Hopf19]). However, Hopf’s functional equation proved intractable, leaving the nature of the attractor unclear for decades. Now, with the loop equation replacing the Hopf functional equation, we have its solution —the Euler ensemble —as the realization of Hopf’s conjecture about the turbulent attractor.
Later, the loop space formalism emerged, rooted in the Makeenko-Migdal (MM) equations of the late 1970s and early 1980s [MMEq79, MM1981NPB, Mig83]. These were immediately recognized in the high-energy physics community as the first exact, non-perturbative dynamical equations for QCD. However, the initial hope of finding an analytical solution for quark confinement soon faded. The mathematical formulation, while geometrically insightful, proved analytically intractable, hampering further progress and shifting the community’s focus toward lattice gauge theory. The loop approach offered a different path compared to Hopf’s formulation, focusing on circulation observables.
The scope of the formalism was extended to fluid dynamics in the early nineties with the derivation of the loop equations for the Navier-Stokes system [M93]. Within this framework, a WKB solution was derived for the limit of large, smooth loops. This solution established a theoretical "area law" for the statistics of velocity circulation. A transformation to a dual "momentum loop space" was also shown to reduce the problem to a singular, one-dimensional equation. For decades, however, further progress was impeded by the mathematical challenges posed by this equation and by a lack of high-precision data with which to test the theory’s predictions.
The theoretical landscape evolved considerably after 2019, when high-resolution direct numerical simulations (DNS) by Sreenivasan and collaborators reported results consistent with the area law derived nearly three decades earlier [S21, Pumir2021Minimal]. This numerical evidence renewed interest in the approach. At the phenomenological level, the statistics of velocity circulation, the Area law, and its various applications to classical and quantum turbulence were discussed in a series of important papers, elucidating both the microscopic meaning, numerical verification by DNS and the applicability of the circulation statistics to the turbulence problem [Polanco2021NatComm, Muller2021PRX, Moriconi2022PRE, Zhu2023PRL, Iyer2024PhysFluids, Moriconi2025arXiv].
This development also provided the impetus to resolve the long-standing mathematical challenges of the momentum loop equation. The complete analytical solution that emerged, realizing aspects of Hopf’s original vision via loop space, is the subject of this review. (See also [migdal2024quantum] for further discussion on the connection to Hopf’s work).
@@
This theoretical structure, initially derived from physical principles, has recently received significant independent validation. From the number theory perspective, the statistical properties of the Euler ensemble at finite $N$ were studied in [Zah23] by combinatorial methods, confirming and extending the results and conjectures made in the first paper [migdal2023exact]. In addition, the distribution of radii of regular star polygons with unit side lengths was recently reproduced by Debmalya Basak using rigorous methods of number theory (private communication, to be published). From a Navier-Stokes dynamics perspective, a rigorous analysis by Brue and De Lellis [DeLellisInprep] has confirmed that the Euler ensemble is an exact solution to the discrete version of the loop equation (Theorem 10.6 in their paper). While the final step—proving the existence of the continuum limit of the fluid dynamics observables as the number of loop segments $N\to\infty$—remains an open mathematical problem, their result provides a firm mathematical foundation for the core of the theory. It confirms that the Euler ensemble is not an approximation but an exact solution at the level of the regularized theory.
Remark 1.
Two other issues were left unresolved in that paper :(i) Proving the global uniqueness of the probability measure on the (loop) space of trajectories for appropriately chosen ensemble of initial conditions, and (ii) The boundedness of error arising from the discretization of the loop equation with or without the “liquid loop approach”. Global uniqueness may not be even true: the fixed points of nonlinear systems like Navier-Stokes equations, as we know from the general RG theory of critical phenomena, are not unique, so there could be other fixed points in addition to the Euler ensemble. The empirical evidence from real and numerical experiments (below) speaks in favor of such a universality of the Euler ensemble. The boundedness of error, in our opinion, is just a temporary issue, soon to be resolved. We developed another approach to the polygonal estimates by Bt]rue and de Lellis, which we plan to develop as a rigorous theorem in collaboration with Google AI lab.
@@
This mathematical validation is now strongly complemented by new high-resolution $4096^{3}$ DNS results from Sreenivasan and collaborators [SreeniAkash2025]. These results verify the theory’s quantitative predictions with high precision, representing a significant development. As discussed in Sections 4.1 and 8, the DNS data for key observables—such as the $E\propto t^{-5/4}$ decay law and the universal, nonlinear effective index—are in excellent, parameter-free agreement with the theoretical curves derived from the Euler ensemble (see Fig. 2).
Remark 2.
Initial results from [SreeniAkash2025] seemed to contradict predictions of the Euler ensemble, but this confusion was cleared in discussions with these authors and is reflected both in their final paper , submitted to this issue, and in this paper below. The issue of apparent violation of universality was that unlike the effective scaling index $\xi_{2}(r,t)$, which is dominated by the bulk part of the decaying energy spectrum, the decay of total energy is sensitive to the small wave-vectors, related to the large scale boundary conditions, which are clearly non-universal. The Euler ensemble was never intended to describe such boundary effects; it was built as a theory of homogeneous turbulence in infinite space. In the particular case of the $4096^{3}$ DNS, these large scale boundary effects are polluted by lattice artifacts ( the integer wave-vectors $k=0,1,2$... in inverse lattice units). The authors of [SreeniAkash2025] investigated the influence of these artifacts by cutting out several first integer values of $k$ in the energy spectrum. The total energy was indeed influenced, and that explained most of the non-universality they observed in energy decay, fitted by K41 scaling laws. The enstrophy is less sensitive to these artifacts, and it matches the Euler ensemble with better accuracy. Finally, there are non-universality corrections from subleading power decay terms, as predicted by the Euler ensemble. These corrections may lead to systematic errors $\penalty 10000\ 10\%$ on top of the lattice artifacts.
Together, these mathematical and numerical results provide strong support for the loop space calculus in the Navier-Stokes equation, confirming the Euler ensemble as a robust solution for decaying turbulence.
@@
This review is organized as follows. The discussion begins with an introduction to the general loop space approach, applicable to both abelian fields in fluid dynamics and nonabelian fields in Yang-Mills theory. The approach is then applied to a hierarchy of turbulence problems, beginning with the foundational case of decaying hydrodynamic turbulence, then the transport of passive scalars, and finally the more complex case of magnetohydrodynamic (MHD) turbulence. A subsequent section situates these analytical solutions within the field’s historical context and clarifies their relationship to classical paradigms of turbulence theory. A particularly striking consequence of this framework emerges in gauge theory. In the Yang–Mills gradient flow, the same loop-space formalism leads to an exact analytic solution—the Hodge-dual matrix surface—which solves the fixed-point Yang–Mills loop equation by harmonic map [migdal2025HodgeDualMatrix]. This result establishes a concrete geometric foundation for the QCD confining string and provides a direct analytical bridge between turbulence, gauge theory, and string dualities.
@@ 2 The Loop Space Approach: From Nonlinear PDEs to Linear Diffusion
Remark 3.
The boundary conditions are vanishing velocity at infinity (or arbitrary constant velocity, which is equivalent by virtue of Galilean invariance). The constant velocity ( or any potential velocity $v_{\mu}=\partial_{\mu}\phi$ ) drops from the circulation we are investigating in this paper.
@@
Remark 4.
The loop $C$ in our functional needs to be continuous for gauge invariance in the Yang-Mills case and for the independence of the potential part of velocity in the Navier-Stokes case. No smoothness is assumed nor needed: the polygonal loop is a valid example. Furthermore, the Functional Fourier transform to the momentum loops would require path integration over coordinate loops $C$, which would make the relevant loops Brownian paths rather than smooth loops. The existence of the dot functional derivatives $\frac{\delta}{\delta\dot{C}(\theta)}$ will be required, but these derivatives exist for finite vorticity regardless of the smoothness of the loop itself. This is based on a Stokes theorem, valid for any continuous loop bounding a surface.
@@
Here, the gauge field $A_{\mu}$ belongs to some Lie algebra, $A_{\mu}=\mathrm{i}\sum T_{a}A^{a}_{\mu}$, which we do not need to specify. The equation for the trace of the ordered path integral (with $\mathbb{P}$ denoting a standard path ordering operator in QFT textbooks)
@@
The transformation of both equations to the loop space diffusion equation is based on the operator calculus invented by Feynman [Feynman1951]. As applied to our case of the loop functionals for fluid dynamics and the gauge theory, it allows us to write both of them as the path ordered exponential times the unit operator
@@
This fundamental geometric identity, proven in the [migdal2025SQYMflow] and our Appendix A, is identical in abelian and non-abelian cases, except that a group space trace is taken in the Yang-Mills case. In both cases, it reveals the geometric meaning of the loop functional as a parallel transport of the covariant derivative operator around a loop $C$, reducing to a c-number for a closed loop. The reduction of the ordered exponential to a c-number is also supported by the Magnus expansion [Magnus1954].
Remark 5.
The term c-number in this context means any operator commuting with all components of covariant derivative operator $D_{\alpha}(x_{0})$, in particular, it could be an arbitrary complex number. This distinction between conventional definition of parallel transport and this operator definition is subtle but critical. In much of the mathematical literature, such as [Gockeler1987], the Wilson loop (the L.H.S. of our identity) is taken as the definition of parallel transport. Our proof, however, establishes this correspondence as a rigorous equality derived from the more fundamental, anchored operator (the R.H.S.). This approach mirrors a historical pattern in theoretical physics where a concept, once viewed as a powerful analogy, is later shown to be a provable, first-principles identity.
@@
Remark 6.
The notion of the diffusion equation may be confusing, as until this moment we were studying a simple phase factor (exponential of circulation for a given velocity field around a fixed or moving loop). The diffusion implies a statistical process with some probability distribution of random variables. How did we switch from a particular solution of the Navier-Stokes (or Yang-Mills) equations to the diffusion process.? Where is the origin of randomness here? This fundamental approach was pioneered by Eberhard Hopf in early fifties [Hopf1952], and it assumes that initial data for velocity field was a statistical distribution ta]rather than a smooth initial field assumed by naive mathematicians as a Cauchy data. In the physical world, water only flows above zero degrees Celsius (ice does not flow, let alone turbulence). Finite temperature means Gibbs statistical ensemble for initial velocity field. These thermal fluctuations, though smaller than the turbulent ones, can trigger the spontaneous stochasticity of the flow [Bernard1998JStatPhys, Thalabard2020CommPhys]. The argumentation of Hopf was more abstract, but it was the same idea of statistical distribution of velocity fields, inherited from initial conditions and evolving by the Navier-Stokes equations. Furthermore, he expressed the idea of a turbulent attractor, which was a prophetic vision of modern theory of turbulence, with the Euler ensemble being precisely the loop scape implementation of Hopf’s turbulent attractor. Technically, the loop equation being a linear functional differential equation, obeys the superposition principle. Therefore, being derived for an arbitrary solution of the Navier-Stokes equation, it also applies to any linear superposition, in particular, to the statistical average of the loop functional over initial Gibbs distribution $\exp\left(-\beta\int v^{2}\right)$ at $t=0$.
Diffusion or quantum mechanics?
The loop equation is formally real, yet the loop operator $\mathcal{L}_{C}$ is a third-order anti-Hermitian operator. Consequently, none of the standard properties of elliptic operators associated with conventional diffusion apply to this loop space dynamics.
In fact, this system is more accurately interpreted as quantum mechanics governed by a Schrödinger equation:
$\mathrm{i}\nu\partial_{t}\tilde{\Psi}[C,t]=\mathcal{H}_{C}\tilde{\Psi}[C,t];$ (22)
$\mathcal{H}_{C}=\mathrm{i}\nu^{2}\mathcal{L}_{C}$ (23)
The viscosity $\nu$ possesses the same physical dimensions $L]^{2}/[T]$ as the kinematic action (analogous to Planck’s constant $\hbar$ per unit mass), and it plays a strictly analogous role in this loop space dynamics. The Hamiltonian is linear in the canonical coordinates $C(\theta)$ and cubic in the canonical momenta $\delta/\delta C(\theta)$, scaling as $C^{-2}$. This resembles the kinetic energy of a free particle, but without the requirement of positivity.
There are zero modes, which are responsible for the fixed points in the Yang-Mills case. However, in the Navier-Stokes case, there is only a trivial zero mode $\Psi[C]=1$.
Crucially, the wave function is complex and exhibits oscillations, a hallmark of quantum mechanics. These oscillations manifest in observable correlation functions as genuine interference effects within this loop space quantum mechanics. It is important to clarify that these quantum features are intrinsic to the continuum loop dynamics and do not arise from the discrete number-theoretic structure of the regular star polygons.
The turbulent limit $\nu\to 0$ corresponds to the WKB limit of this loop space quantum mechanics. This analogy is the key to the analytic solution of the turbulence problem described in this review.
In the remainder of this paper, we shall refer to this dynamics as "loop space diffusion" to maintain familiarity for practitioners of fluid dynamics. However, make no mistake: this is an imaginary diffusion, better known as quantum mechanics.
This raises an intriguing philosophical question: how can a deterministic classical system exhibit quantum interference effects? First, the system is not truly deterministic in the reversible sense: the imaginary part of the loop functional $\Psi[C]$ is directly related to energy dissipation, leading to a violation of time-reversal symmetry. The probabilistic distribution of dissipative turbulent flow possesses all the features of a quantum mechanical wave function, including the quantum interference of amplitudes in the sum over histories. The Schrödinger equation here is not a poetic analogy but an exact mathematical equivalence, much like the thermal radiation of black holes.
Second, while complex numbers may appear to be a mathematical artifice, they are as fundamental here as the complex amplitudes and phases of electromagnetic waves in classical electrodynamics. The phenomenon of wave interference is an objective physical process, and complex amplitudes are the necessary mathematical tools to describe it. The same logic applies to the circulation dynamics in turbulent flow. As we shall see below, these quantum oscillations of the loop amplitudes manifest as experimentally observed oscillations in the energy spectrum of turbulent flow, such as those recorded in the wind tunnel at the Max Planck Institute Turbulence Lab. The interference is real; its mathematical mechanism is identical to that of quantum mechanics. Whether one labels it "quantum" or "dissipative statistical" is merely a matter of semantics.
@@ 3 Momentum loop equation
Remark 7.
Let us clarify this important point. This equation for $P(\theta,\tau)$ is an ordinary differential equation with time derivative $\partial_{\tau}P(\theta,\tau)$ being a nonlinear functional of $P$ depending on infinitesimal neighbors on a unit circle $P(\theta\pm 0,\tau)$. This equation is singular but deterministic. The randomness comes from the initial data $P(\theta,0)$, which are random variables with the distribution given by Fourier transform of the initial $\Psi(C)$. This distribution was computed in [migdal2025duality], but we do not need it here. The idea behind this MLE is that this solution $P(\theta,\tau)$ starts from a specific random data $P(\theta,0)$, determined by transforming the initial Gibbs average $\left\langle\Psi(C)\right\rangle_{\mathrm{Gibbs}}$ into momentum loop space, but in course of evolution, it is attracted by a fixed point (the Hopf’s turbulent attractor). This fixed point is a degenerate solution of the MLE, depending upon infinite number of arbitrary parameters. In other words, this fixed point is a set or a manifold, like the Gibbs distribution was a manifold (the sphere of constant energy in infinite dimensional space of velocity field). Where is this attractor located? Not in the physical space, like inside the bulk of the flow. Not even in the loop space $C(\theta)$– it is located in the momentum loop space $P(\theta)$. The momenta are related to coordinates as hamiltonian conjugates, which means that discrete lov=calized structures in momentum space correspond to condtinuous distributions in coordinate space, – here in the loop space. So, the regular star polygons of the Euler ensemble sit in abstract space, not reflecting in the discrete structures in pg=hysical space (with a notable exception of passive scalar, we consider later in this review). Given the attractor maniofold, we follow the ergodic theory by assuming that the time evolution uniformly covers this manifold with invariant measure, determined by its internal mathematical symmetries. The miracle is that in the Navier-Stokes MLE, this manifold can be analytically found, its invariant measure computed using modern Number Theory, and the emerging turbulent distribution investigated in the local limit, when it also has infinitely many degrees of freedom. The idea of a fixed point is a basis of Wilson’s RG approach to critical phenomena, where it is more a philosophy than an analytical method. Not so in the turbulence problem! Here we have a priceless gift of analytic solution rooted in the Number Theory.
@@
The relation between the loop operator and the discontinuity of the momentum loop $\Delta P_{\nu}(\theta)$ was discovered and investigated in earlier papers in QCD [MLDMig86, SecQuanM95, Mig98Hidden]. The Momentum loop equation for QCD involves contact terms that provide boundary conditions in loop space for self-intersecting loops. We are not studying these terms in the present paper, restricting ourselves to the Yang-Mills gradient flow. For QCD, this equation corresponds to the WKB limit of large loops without self-intersections, when some form of the area law is expected as an asymptotic solution.